### v1

@@ 3 Preliminaries
One of the main ingredients in our proof is a variant of (classical) the relational adversary method from quantum computing given in [BCR24].
@@
Consider finite sets $A$ and $B$, a set $\mathcal{X}\subseteq B^{A}$ of functions, and a map $\mathcal{H}:\mathcal{X}\to\{0,1\}$ which assigns a label to each function in $\mathcal{X}$. Additionally, we get oracle access to an unknown function $F^{*}\in\mathcal{X}$. The problem is to compute $\mathcal{H}(F^{*})$ using as few queries to $F^{*}$ as possible. [^8] In other words, we have free access to $\mathcal{H}$ and the only queries counted are the ones to $F^{*}$, which will be of the form: “What is $F^{*}(a)$?”, for some $a\in A$. The oracle will return $F^{*}(a)$ in one computational step. Let $r:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ be a non-zero symmetric function of our choice with $r(F_{1},F_{2})=0$ whenever $\mathcal{H}(F_{1})=\mathcal{H}(F_{2})$. For each $\mathcal{Z}\subseteq\mathcal{X}$, define
@@ 4.1 Setup
Having defined the relation $r$, applying Lemma 1 will give a lower bound by considering a distribution $p$ over $\mathcal{X}$, where each function $F\in\mathcal{X}$ is given as input with probability $p(F)=\frac{M(\{F\})}{M(\mathcal{X})},$ where $M$ is as defined in Eq. 4 for the relation $r$. What remains to be done is to explain how to invoke Lemma 1 and estimate the lower bound it gives when the input distribution is $p$.
@@ 4.2 Proof of the main theorem
First, we may assume that $n\geq 16\sigma^{2}$. This is because in the alternative case where $n<16\sigma^{2}$, the theorem statement does not give anything useful as $\sqrt{n}<\exp(3\sigma)$ and $t_{mix}(\sigma/(2n))\geq 1$. The value functions we will use are of the form $f_{\mathbf{x}}$ as seen in Definition 5. These functions are parametrized by walks $\mathbf{x}$ of length $L$ from the set $\mathcal{W}$ defined in 1. For sake of invoking Lemma 1 however, we must turn the local search problem into a decision problem. To do this, we use the technique shown in  [Aar06, DR10]: associate with each function $f_{\mathbf{x}}$ the function $g_{\mathbf{x},b}$ defined in Definition 5. This hides a bit at the local minimum vertex (while hiding the value $-1$ at every other vertex). The decision problem is: given the graph $G$ and oracle access to a function $g_{\mathbf{x},b}$, return the hidden bit $b$; i.e. we set the function $\mathcal{H}$ for use in Lemma 1 as $\mathcal{H}(g_{\mathbf{x},b})=b$. By 2, the function $f_{\mathbf{x}}$ as defined in Definition 5 is valid. Therefore by Lemma 10, $f_{\mathbf{x}}$ has a unique local minimum, namely $x_{L}$. This means that $g_{\mathbf{x},b}$ as defined in Definition 5 does indeed hide the bit $b$ only at the local minimum of $f_{\mathbf{x}}$. Therefore measuring the query complexity of this decision problem will give the answer for local search, as the following two problems have query complexity within additive $1$:
  • search problem: given oracle access to a function $f_{x}$, find a vertex $v$ that is a local minimum;
  • decision problem: given oracle access to a function $g_{x,b}$, find $b$.
We then invoke Lemma 1 with $\mathcal{X}$ as defined in Definition 5 and with $\mathcal{H}(g_{\mathbf{x},b})=b$. This tells us that the randomized query complexity of the decision problem is at least $\min\limits_{\begin{subarray}{c}\mathcal{Z}\subseteq\mathcal{X}:q(\mathcal{Z})>0\end{subarray}}0.01\cdot{M(\mathcal{Z})}/{q(\mathcal{Z})}$, with $M(\mathcal{Z})$ and $q(\mathcal{Z})$ as defined in Lemma 1. By Lemma 7, there exists a subset $\mathcal{Z}\subseteq\mathcal{X}$ with $q(\mathcal{Z})>0$. From this point on, we fix an arbitrary subset $\mathcal{Z}\subseteq\mathcal{X}$ of functions with $q(\mathcal{Z})>0$ and will then lower bound $M(\mathcal{Z})$ and upper bound $q(\mathcal{Z})$.
@@
$\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}$ (85)
@@
where in Eq. 87 we used the inequality $a\leq b+|a-b|$ for $a,b\geq 0$. A random walk with starting vertex drawn from $\pi$ has probability $\pi(v)$ of being at $v$ at each step. Formally, for all $\ell\in\mathbbm{N}$ we have
@@ Appendix A Helper lemmas
$i=j,\ldots,\lfloor\sqrt{n}\rfloor$, sample the segment $Tail(\mathbf{z},i-1,i)$ conditioned on having set $Head(\mathbf{z},i-1)$. For every $0\leq k\leq\lfloor\sqrt{n}\rfloor$, the set of vertices given by the first $k+1$ milestones of $\mathbf{z}$ is
@@
Given that $Head(\mathbf{x},j)=Head(\mathbf{z},j)$, the condition $J(\mathbf{x},\mathbf{z})=j$ is equivalent to
@@
$\qquad=\Pr\Bigl{[}\Bigl{(}z_{kT}\not\in S_{k-1}\;\forall k\in\{j+1,\ldots,\lfloor\sqrt{n}\rfloor\}\Bigr{)}\land\Bigl{(}z_{(j+1)T}\neq x_{(j+1)T}\ \Bigr{)}\mid Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}$ (133)
@@
This concludes the proof of the lemma. ∎ The next lemma is inspired by Lemma 9 from [BCR24]. However, it is slightly different, so we include the proof here.
@@
$\sum_{g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\mathbbm{1}_{\{g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)\}}$ $\leq\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{x},J(\mathbf{x},\mathbf{y}))\cup Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})$ (157)
$\leq\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{x},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})+\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})$ (162)
@@
Combining LABEL:eq:exists_Z_1 and Eq. 168, we get $s\leq n/16$. Thus the number of vertices reachable from $u$ in $T$ steps is $n-s\geq 15n/16$. For $n\geq 5$, we have $215n/16\geq\lfloor\sqrt{n}\rfloor+$ 2, which means that at least $\lfloor\sqrt{n}\rfloor+2$ vertices are reachable via $\mathcal{P}$ from any vertex $u$ within $T$ steps. We can then construct two walks, $\mathbf{x}$ and $\mathbf{y}$, in the following manner. For $i=1,\ldots,\lfloor\sqrt{n}\rfloor$, take $x_{iT}$ to be an arbitrary vertex reachable from $x_{(i-1)T}$ other than $x_{0},\ldots,x_{(i-1)T}$; this is possible since at least $\lfloor\sqrt{n}\rfloor+2$ vertices are reachable. Connect the milestones using an arbitrary path such that every edge $(u,w)$ in the path has $\mathcal{P}_{u,w}>0$. Construct $\mathbf{y}$ in the same manner but requiring that $Head(\mathbf{y},\lfloor\sqrt{n}\rfloor-1)=Head(\mathbf{x},\lfloor\sqrt{n}\rfloor-1)$ and $y_{L}\not\in\left\{x_{0},x_{T},x_{2T},\ldots,x_{\lfloor\sqrt{n}\rfloor T}\right\}$. Define $\mathcal{Z}=\{g_{\mathbf{x},0},g_{\mathbf{y},1}\}$. We will show that $q(\mathcal{Z})>0$. First observe that $\mathbf{x}\neq\mathbf{y}$, both $\mathbf{x}$ and $\mathbf{y}$ are good since they do not repeat milestones, and the bit hidden by $g_{\mathbf{x},0}$ is different from the bit hidden by $g_{\mathbf{y},1}$. The length of the prefix shared by $\mathbf{x}$ and $\mathbf{y}$ is $J(\mathbf{x},\mathbf{y})=\lfloor\sqrt{n}\rfloor-1$. Then
@@ Appendix B Corollaries of the main theorem
We obtain the following corrollary, which lower bounds the randomized complexity of local search as a function of the spectral gap of $\mathcal{P}$. See 1
@@
As a special case, consider the simple lazy random walk on a graph with $d_{max}/d_{min}\leq C$ for constant $C$. Applying Lemma 12 to this walk yields the following result: See 1
@@
Substituting Eq. 181 into Lemma 12, we get a lower bound of $\Omega\bigl{(}\frac{(1-\lambda_{2})\sqrt{n}}{\log^{2}(n)}\bigr{)}$ on the randomized query complexity of local search on $G$. This completes the proof of the corollary. ∎ Compare to the following corollary, which is just 1 applied to the simple lazy random walk when $d_{max}/d_{min}$ is bounded by a constant $C$. See 2

### v3

@@ 3 Preliminaries
One of the main ingredients in our proof is a variant of the (classical) relational adversary method from quantum computing given in [BCR24].
@@
Consider finite sets $A$ and $B$, a set $\mathcal{X}\subseteq B^{A}$ of functions, and a map $\mathcal{H}:\mathcal{X}\to\{0,1\}$ which assigns a label to each function in $\mathcal{X}$. Additionally, we get oracle access to an unknown function $F^{*}\in\mathcal{X}$. The problem is to compute $\mathcal{H}(F^{*})$ using as few queries to $F^{*}$ as possible. [^8] In other words, we have free access to $\mathcal{H}$ and the only queries counted are the ones to $F^{*}$, which will be of the form: “What is $F^{*}(a)$?”, for some $a\in A$. The oracle will return $F^{*}(a)$ in one computational step.
Let $r:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{\geq 0}$ be a non-zero symmetric function of our choice with $r(F_{1},F_{2})=0$ whenever $\mathcal{H}(F_{1})=\mathcal{H}(F_{2})$. For each $\mathcal{Z}\subseteq\mathcal{X}$, define
@@ 4.1 Setup
This value $r(\mathbf{x},\mathbf{y})$ may be interpreted as the probability that two walks sampled from the Markov chain will be $\mathbf{x}$ and $\mathbf{y}$ respectively, conditioned on diverging for the first time at an index in the range $(jT,(j+1)T]$.
@@
Having defined the relation $r$, applying Lemma 1 will give a lower bound by considering a distribution $p$ over $\mathcal{X}$, where each function $F\in\mathcal{X}$ is given as input with probability $p(F)=\frac{M(\{F\})}{M(\mathcal{X})},$ where $M$ is as defined in Eq. 4 for the relation $r$.
What remains to be done is to explain how to invoke Lemma 1 and estimate the lower bound it gives when the input distribution is $p$.
@@ 4.2 Proof of the main theorem
First, we may assume that $n\geq 16\sigma^{2}$. This is because in the alternative case where $n<16\sigma^{2}$, the theorem statement does not give anything useful as $\sqrt{n}<\exp(3\sigma)$ and $t_{mix}(\sigma/(2n))\geq 1$.
The value functions we will use are of the form $f_{\mathbf{x}}$ as seen in Definition 5. These functions are parametrized by walks $\mathbf{x}$ of length $L$ from the set $\mathcal{W}$ defined in 1. For sake of invoking Lemma 1 however, we must turn the local search problem into a decision problem. To do this, we use the technique shown in  [Aar06, DR10]: associate with each function $f_{\mathbf{x}}$ the function $g_{\mathbf{x},b}$ defined in Definition 5. This hides a bit at the local minimum vertex (while hiding the value $-1$ at every other vertex). The decision problem is: given the graph $G$ and oracle access to a function $g_{\mathbf{x},b}$, return the hidden bit $b$; i.e. we set the function $\mathcal{H}$ for use in Lemma 1 as $\mathcal{H}(g_{\mathbf{x},b})=b$.
By 2, the function $f_{\mathbf{x}}$ as defined in Definition 5 is valid. Therefore by Lemma 10, $f_{\mathbf{x}}$ has a unique local minimum, namely $x_{L}$. This means that $g_{\mathbf{x},b}$ as defined in Definition 5 does indeed hide the bit $b$ only at the local minimum of $f_{\mathbf{x}}$. Therefore measuring the query complexity of this decision problem will give the answer for local search, as the following two problems have query complexity within additive $1$:
  • search problem: given oracle access to a function $f_{x}$, return a vertex $v$ that is a local minimum;
  • decision problem: given oracle access to a function $g_{x,b}$, return $b$.
We then invoke Lemma 1 with $\mathcal{X}$ as defined in Definition 5 and with $\mathcal{H}(g_{\mathbf{x},b})=b$. This tells us that the randomized query complexity of the decision problem is at least $\min\limits_{\begin{subarray}{c}\mathcal{Z}\subseteq\mathcal{X}:q(\mathcal{Z})>0\end{subarray}}0.01\cdot{M(\mathcal{Z})}/{q(\mathcal{Z})}$, with $M(\mathcal{Z})$ and $q(\mathcal{Z})$ as defined in Lemma 1.
By Lemma 7, there exists a subset $\mathcal{Z}\subseteq\mathcal{X}$ with $q(\mathcal{Z})>0$. From this point on, we fix an arbitrary subset $\mathcal{Z}\subseteq\mathcal{X}$ of functions with $q(\mathcal{Z})>0$ and will then lower bound $M(\mathcal{Z})$ and upper bound $q(\mathcal{Z})$.
@@
$\sum_{j=0}^{\lfloor\sqrt{n}\rfloor-1}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ v\in Tail(\mathbf{y},j+1)\\ J(\mathbf{x},\mathbf{y})\geq j\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}$
@@
where in Eq. 29 we used the inequality $a\leq b+|a-b|$ for $a,b\geq 0$.
A random walk with starting vertex drawn from $\pi$ has probability $\pi(v)$ of being at $v$ at each step. Formally, for all $\ell\in\mathbbm{N}$ we have
@@
Acknowledgements
We would like to thank Davin Choo for several useful discussions.
@@ Appendix A Helper lemmas
$i=j,\ldots,\lfloor\sqrt{n}\rfloor$, sample the segment $Tail(\mathbf{z},i-1,i)$ conditioned on having set $Head(\mathbf{z},i-1)$.
For every $0\leq k\leq\lfloor\sqrt{n}\rfloor$, the set of vertices given by the first $k+1$ milestones of $\mathbf{z}$ is
@@
Given that $Head(\mathbf{x},j)=Head(\mathbf{z},j)$, a sufficient condition for $J(\mathbf{x},\mathbf{z})=j$ is
@@
$\qquad\geq\Pr\Bigl{[}\Bigl{(}z_{kT}\not\in S_{k-1}\;\forall k\in\{j+1,\ldots,\lfloor\sqrt{n}\rfloor\}\Bigr{)}\land\Bigl{(}z_{(j+1)T}\neq x_{(j+1)T}\ \Bigr{)}\mid Head(\mathbf{x},j)=Head(\mathbf{z},j)\Bigr{]}$ (53)
@@
$\sum_{\begin{subarray}{c}\mathbf{y}\in\mathcal{W}:\\ J(\mathbf{x},\mathbf{y})=j\\ good(\mathbf{y})\end{subarray}}\mathcal{P}[Tail(\mathbf{y},j)]\cdot\mathcal{P}_{y_{j\cdot T},\;y_{j\cdot T+1}}\geq 2^{-4\sigma}\,.$
This concludes the proof of the lemma. ∎
The next lemma is inspired by Lemma 9 from [BCR24]. However, it is slightly different, so we include the proof here.
@@
$\sum_{g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})\mathbbm{1}_{\{g_{\mathbf{x},b_{1}}(v)\neq g_{\mathbf{y},b_{2}}(v)\}}$ $\leq\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{x},J(\mathbf{x},\mathbf{y}))\cup Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})$
$\leq\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{x},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})+\sum_{\begin{subarray}{c}g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}}\in\mathcal{Z}:\\ v\in Tail(\mathbf{y},J(\mathbf{x},\mathbf{y}))\end{subarray}}r(g_{\mathbf{x},b_{1}},g_{\mathbf{y},b_{2}})$
@@
Combining LABEL:eq:exists_Z_1 and Eq. 74, we get $s\leq n/16$. Thus the number of vertices reachable from $u$ in $T$ steps is $n-s\geq 15n/16$. For $n\geq 5$, we have $215n/16\geq\lfloor\sqrt{n}\rfloor+$ 2, which means that at least $\lfloor\sqrt{n}\rfloor+2$ vertices are reachable via $\mathcal{P}$ from any vertex $u$ within $T$ steps.
We can then construct two walks, $\mathbf{x}$ and $\mathbf{y}$, in the following manner. For $i=1,\ldots,\lfloor\sqrt{n}\rfloor$, take $x_{iT}$ to be an arbitrary vertex reachable from $x_{(i-1)T}$ other than $x_{0},\ldots,x_{(i-1)T}$; this is possible since at least $\lfloor\sqrt{n}\rfloor+2$ vertices are reachable. Connect the milestones using an arbitrary path such that every edge $(u,w)$ in the path has $\mathcal{P}_{u,w}>0$. Construct $\mathbf{y}$ in the same manner but requiring that $Head(\mathbf{y},\lfloor\sqrt{n}\rfloor-1)=Head(\mathbf{x},\lfloor\sqrt{n}\rfloor-1)$ and $y_{L}\not\in\left\{x_{0},x_{T},x_{2T},\ldots,x_{\lfloor\sqrt{n}\rfloor T}\right\}$.
Define $\mathcal{Z}=\{g_{\mathbf{x},0},g_{\mathbf{y},1}\}$. We will show that $q(\mathcal{Z})>0$. First observe that $\mathbf{x}\neq\mathbf{y}$, both $\mathbf{x}$ and $\mathbf{y}$ are good since they do not repeat milestones, and the bit hidden by $g_{\mathbf{x},0}$ is different from the bit hidden by $g_{\mathbf{y},1}$. The length of the prefix shared by $\mathbf{x}$ and $\mathbf{y}$ is $J(\mathbf{x},\mathbf{y})=\lfloor\sqrt{n}\rfloor-1$. Then
@@ Appendix B Corollaries of the main theorem
We obtain the following corrollary, which lower bounds the randomized complexity of local search as a function of the spectral gap of $\mathcal{P}$.
See 1
@@
As a special case, consider the simple lazy random walk on a graph with $d_{max}/d_{min}\leq C$ for constant $C$. Applying Lemma 12 to this walk yields the following result:
See 1
@@
Substituting Eq. 87 into Lemma 12, we get a lower bound of $\Omega\bigl{(}\frac{(1-\lambda_{2})\sqrt{n}}{\log^{2}(n)}\bigr{)}$ on the randomized query complexity of local search on $G$. This completes the proof of the corollary. ∎
Compare to the following corollary, which is just 1 applied to the simple lazy random walk when $d_{max}/d_{min}$ is bounded by a constant $C$.
See 2